one of the zeros of the function f(x)=x^3+5x^2-9x-45 is x=-5, find the zeros of the function
Answer by lwsshak3(11628)

one of the zeros of the function f(x)=x^3+5x^2-9x-45 is x=-5, find the zeros of the function
To find the other zeros or roots of the given third degree equation,
divide the given equation by x+5 by long division method or if you know how to do synthetic division, use -5 as the root.
With this format I am working with, I cannot show you how to do this, but if you do it correctly, you will get a reduced second degree equation, x^2-9, which is a difference of two squares.
In factored form f(x) = (x+5)(x^2-9)=(x+5)(x+3)(x-3)
The zeros or roots of f(x)are:-5,-3,and 3